OpenStudy (anonymous):

Exact value of sin 7 pi?

OpenStudy (anonymous):

I just figured it out nevermind ^^

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm sin(7π)=sin(6π+π)=sin(π) }\)

OpenStudy (solomonzelman):

This is the simplest way. You could have converted radians to degrees and use sin(a+b). Which is the long version of what I did.